OpenStudy (anonymous):

Solve the following equation analytically. 6e^2x = 24e^−6x

OpenStudy (anonymous):

Suggestion: Take the natural logarithm of each side of the equation.

OpenStudy (anonymous):

how would i begin that?

OpenStudy (anonymous):

\[\ln 6e ^{2x} = \ln 6 + 2x and \ln 24e ^{-6x} = \ln 24 - 6x\] You now have the more familiar ln 6 + 2x = ln 24 - 6x. ln 6 might look horrible but it's just a constant term.